Science Report – April 29th

Science Report for 29 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no sites were noted as useful for emergency evacuation shelters. If a crew was trapped north on Copernicus highway, there are several low sites where they could shelter out of the wind.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found two dark features on the sun and another possible solar flare.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: The first EVA of the day took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Final samples of the project were taken from a towel wiped dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria growth and monitoring will continue over the remaining days of the mission.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No fogging was reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Crew 193 XO and Science Officer

Copyright © The Mars Society. All rights reserved. | Main Site